An 18 quart radiator contains a 40% solution of antifreeze. How much must be drained out and replaced with pure antifreze to obtain a 70% antifreeze solution?
Answer by checkley77(12569)

.40(18-x)+x=.7*18
7.2-.4x+x=12.6
.6x=12.6-7.2
.6x=5.4
x=5.4/.6
x=9 qts. of pure antifreeze is used.
Proof:
.40(18-9)+9=.7*18
.40*9+9=12.6
3.6+9=12.6
12.6=12.6